A 30-kg rock is on top of a 50-m cliff. How much of its potential energy is lost when it falls and
reaches 20 m above the ground
1
Expert's answer
2020-11-23T10:28:24-0500

The loss om potential energy is:


"\Delta U = |U_2 - U_1|"

Where "U_1" is the initial energy and "U_2" is the final potential energy of the rock. The initial is:


"U_1 = mgh_1"

where "m = 30kg" is the mass of the rock, "g = 9.81m\/s^2" is the gravitational acceleration, and "h_1 = 50m" is the initial height. Thus, obtain:


"U_1 = 30\cdot 9.81\cdot 50 = 14715J"

Similarly, the final potential energy is:


"U_2 = mgh_2"

where "h_2 = 20m" is the final height of the rock. Thus, get:


"U_2 = 30\cdot 9.81\cdot 20 = 5886J"

Then the loss is:


"\Delta U = |5886J-14715J| = 8829J"

Answer. 8829 J.